OpenStudy (anonymous):

use quotient rule to simplify: √(7/10)

OpenStudy (anonymous):

do you mean rationalize the denominator?

OpenStudy (anonymous):

?? sry

OpenStudy (anonymous):

what does it mean "simplify"?

OpenStudy (anonymous):

sqrt of 7 divide 10 i think

OpenStudy (anonymous):

maybe you are supposed to day \[\sqrt{\frac{7}{10}}=\frac{\sqrt{7}}{\sqrt{10}}\] \[=\frac{\sqrt{7}}{\sqrt{10}}\times \frac{\sqrt{10}}{\sqrt{10}}\] \[=\frac{\sqrt{70}}{10}\]

OpenStudy (anonymous):

That looks right, but the answer wasn't what it was looking for. Says "Recall that the square foot of a quotient is the quotient of the square roots."

OpenStudy (anonymous):

yes that was step one

OpenStudy (anonymous):

oh, lol ;)

OpenStudy (anonymous):

\[\sqrt{\frac{7}{10}}=\frac{\sqrt{7}}{\sqrt{10}}\] square root of quotient on the left, quotient of square roots on the rigth

OpenStudy (anonymous):

forgive me... I'm an english student stuck in an algebra class!!

OpenStudy (anonymous):

that sucks

OpenStudy (anonymous):

but at least if you know english you understand "square root of quotient" and "quotient of square root" right?

OpenStudy (anonymous):

right, thank you